OpenStudy (anonymous):

If tanθ = 1/3 , then secθ = ?

OpenStudy (anonymous):

\[\sin^2x +\cos^2x=1.\]

OpenStudy (solomonzelman):

Tangent = Opposite / Adjacent (So the missing side is the Hypotenuse, opposite=1, adjacent=3) You know that: (Opposite Side)\(\large\color{black}{ ^2 }\)+(Adjacent Side)\(\large\color{black}{ ^2 }\)=(Hypotenuse)\(\large\color{black}{ ^2 }\) Now, plug in your values \(\large\color{black}{ 1^2+3^2=x^2 }\) (Labeling the hypotenuse with 'x' since it is the unknown)

OpenStudy (solomonzelman):

Can you find the hypotenuse for me please?

OpenStudy (anonymous):

would it be 100?

OpenStudy (solomonzelman):

I don't know how you get 100 for the hypotenuse.